OpenStudy (anonymous):

tan inverse cos x/1-sin x write in simplest form?

OpenStudy (anonymous):

What's wrong with your picture?

OpenStudy (anonymous):

ans will be \[\pi/4 + x/2\]

OpenStudy (anonymous):

do you mean \[\frac{\tan(\cos^{-1}x)}{1-\sin x}\]?

OpenStudy (anonymous):

hoe to get there/

OpenStudy (anonymous):

after tan inverse i+tanx\2 /1- tan x\2 wt is the next step/

OpenStudy (anonymous):

s to lifesaver

OpenStudy (anonymous):

I meant... did you ask to simplify the expression I wrote down?

OpenStudy (anonymous):

s sir

OpenStudy (anonymous):

step for solving 1- conver cosx into cos^x/2- sin^2x/2 and 1- sinx = sin^2x/2 + Cos^2x/2 - 2sinx/2cosx/2

OpenStudy (anonymous):

can u say step by step plz

OpenStudy (anonymous):

Let x = cos y. Then \[\sqrt{1-x^2}=\sqrt{1-\cos^2y}=\sqrt{\sin^2y}=\sin y\] Hence \[\tan(\cos^{-1}x)=\tan y=\frac{\sin y}{\cos y}=\frac{\sqrt{1-x^2}}{x}\] and therefore \[\frac{\tan(\cos^{-1}x)}{1-\sin x}=\frac{\sqrt{1-x^2}}{x(1-\sin x)}\] I think this doesn't simplify further.

OpenStudy (anonymous):

u can call me and ask this problem... my no is +91 -9706994872

OpenStudy (anonymous):

sir i don't have a mob

OpenStudy (anonymous):

does this ans satissfies my above question?

OpenStudy (anonymous):

yes this is the only ans for ur question.... π/4+x/2

OpenStudy (anonymous):

how this both links sir?

OpenStudy (anonymous):

step for solving 1- conver cosx into cos^x/2- sin^2x/2 and 1- sinx = sin^2x/2 + Cos^2x/2 - 2sinx/2cosx/2 after this just factorize Nr and Dr and then ...it becomes tan(pi/4+x) so tan inverse cancil out tan and ans will be π/4+x/2

OpenStudy (anonymous):

if i dont disturb u can u wright it in steps i can't read it plz

OpenStudy (anonymous):

hey r u fed up with me/

OpenStudy (anonymous):

ok u will send me sir i got to go for bath.